simplify:
(2x^3+3x^2+2x+2)+(x+1)
I have 4 choices for answer, but my work never gets close to any of them:
a) 2x^2+3x+1 r2
b) 2x^2+3x+1
c) 2x^2+x+1
d) 2x^2+x+1 r1
I really need to see clear instructions on how to compute this type of problem. Thank you very much
: simplify:
(2x^3+3x^2+2x+2)+(x+1)
I have 4 choices for answer, but my work never gets close to any of them:
a) 2x^2+3x+1 r2
b) 2x^2+3x+1
c) 2x^2+x+1
d) 2x^2+x+1 r1
I really need to see clear instructions on how to compute this type of problem. Thank you very much

Answer by oscargut(666)

simplify:
(2x^3+3x^2+2x+2)+(x+1)
I think the correct question is:
(2x^3+3x^2+2x+2)/(x+1)
doing synthetic division we have
2x^3+3x^2+2x+2=(x+1)(2x^2+x+1)+1 = (x+1)Q(x)+r
so correct option is d)